6a=2-b use the method of substiution
6a=6-b

Found 2 solutions by lynnlo, Ave:
Answer by lynnlo(4176)
Answer by Ave(106)

pick one equation and solve for a or b
6a=6-b divide by 6
a=1-b/6
now that you know what a is, plug in all that into the other equation where a is
6(1-b/6)=2-b
6-b=2-b add b to both sides
6=2 is false, these lines are parallel so the answer is empty set, or no solution